# Automated Self-Testing as a Quality Gate: Evidence-Driven Release Management for LLM Applications

### 13 Mar 2026 v1

@@ Abstract
LLM applications are AI systems whose non-deterministic outputs and evolving model behavior make traditional testing insufficient for release governance. We present an automated self-testing framework that introduces quality gates with evidence-based release decisions (PROMOTE/HOLD/ROLLBACK) across five empirically grounded dimensions: task success rate, research context preservation, P95 latency, safety pass rate, and evidence coverage. We evaluate the framework through a longitudinal case study of an internally deployed multi-agent conversational AI system with specific marketing capabilities in active development, covering 38 evaluation runs across 20+ internal releases. The gate identified two ROLLBACK-grade builds in early runs and supported stable quality evolution over a four-week staging lifecycle while exercising persona-grounded, multi-turn, adversarial, and evidence-required scenarios. Statistical analysis (Mann-Kendall trends, Spearman correlations, bootstrap confidence intervals), gate ablation, and overhead scaling indicate that evidence coverage is the primary severe-regression discriminator and that runtime scales predictably with suite size. A human calibration study ($n=60$ stratified cases, two independent evaluators, LLM-as-judge cross-validation) reveals complementary multi-modal coverage: LLM-judge disagreements with the system gate ($\kappa=0.13$) are attributable to structural failure modes—latency violations and routing errors—invisible in response text alone, while the judge independently surfaces content quality failures missed by structural checks, validating the multi-dimensional gate design. The framework, supplementary pseudocode, and calibration artifacts are provided to support AI-system quality assurance and independent replication.
@@ 1 Introduction
LLM-based applications are rapidly expanding across the software industry, especially in domains such as marketing automation and conversational analytics. Yet teams still rely on QA practices designed for deterministic software, and that mismatch creates severe limitations. Unlike traditional programs, LLM agents are probabilistic, prone to “hallucinations”, and heavily context-dependent [6]. As a result, standard unit and integration tests often miss complex behavioral regressions (e.g., persona drift, subtle safety violations, or unsupported evidence claims).
This mismatch creates a persistent release question for engineering teams: when is a new conversational AI release ready for production? Beyond functional correctness, teams must detect subtle persona drift, enforce guardrails consistently, and prevent regressions in reasoning behavior [6]. The current ecosystem still lacks automated, data-driven end-to-end testing frameworks that evaluate qualitative dimensions from a realistic user-perception perspective, which limits reliable Continuous Deployment in LLM systems.
To mitigate this gap, we present an empirical testing framework based on Automated Self-Testing (historically known as dogfooding [10]), designed to serve as Quality Gates prior to the release of complex LLM systems. By creating a continuous loop where the system is intensively evaluated against a static and dynamic “question bank,” our framework assesses the build across five empirically grounded dimensions—Task Success Rate (build stability and response utility), Research Context Preservation (multi-turn context continuity), P95 Latency (user experience), Safety Pass Rate (guardrail adherence), and Evidence Coverage (factuality anchoring)—selected to comprehensively evaluate both technical robustness and user-perceived quality. Based on these metrics, the framework generates a clear release management decision: PROMOTE/HOLD/ROLLBACK.
We validated our method through a longitudinal empirical study (case study) using an internally deployed multi-agent orchestration system (i.e., a conversational AI agent with specific marketing and analytics capabilities) in active development. We analyzed 38 evaluation runs with question banks evolving from 59 to 133 scenarios across more than 20 internal system releases. The framework correctly identified 2 ROLLBACK-grade builds and supported stable quality evolution over a four-week staging lifecycle. Statistical analysis reveals that evidence coverage was the primary determinant of release rejection, while success rate exhibits a controlled decrease as the question bank grows more challenging.
@@
    A formalized and scalable self-testing framework used as a methodological barrier (quality gate) for Agentic AI.
@@
    A defined and publicly tested set of quality acceptance thresholds for systemic approval (PROMOTE/HOLD/ROLLBACK).
  4. 4.
    Supplementary material, including illustrative examples, pseudocode for the evaluation infrastructure, and sample test cases, providing practical guidance for community adaptation without exposing proprietary assets.
@@ 2.2 Automated Evaluation of LLMs
In contrast to isolated benchmarking, evaluating production systems requires addressing the real-world usage distribution. Our framework builds upon automated evaluation but integrates it directly into the release pipeline as a strict quality gate, rather than an offline benchmarking tool.
@@ 3.2 Quality Dimensions and Thresholds
  • Latency (P95 << 15,000ms): The 95th percentile response time must not exceed 15 seconds, ensuring acceptable real-time user UX despite multi-agent orchestration overhead.
@@
Task Success Rate ($\geq$80%) was set to permit iterative feature development while blocking builds with systemic failures. Pre-study runs established a natural baseline of approximately 92–95%; the 80% threshold provides a 12–15 percentage-point safety margin, absorbing test suite expansion without triggering false rollbacks.
@@
P95 Latency (<<15,000ms) reflects the upper bound of acceptable wait time for an interactive marketing analytics assistant, consistent with general UX guidelines for web applications [16]. Responses exceeding 15 seconds consistently trigger user abandonment in pre-launch internal testing.
@@ 3.4 Decision Engine
This evidence-driven approach ensures that behavioral regressions and subtle degradations—such as an increase in hallucination rates disguised as confident answers—are caught deterministically before impacting users. Figure 3 summarizes the release-decision state machine, and Table 4 positions automated self-testing relative to traditional U/I and E2E testing coverage.
@@ 4.3 Methodology and Data Collection
During this longitudinal period, every code merge targeting the main branch triggered an automated self-testing workflow within an isolated staging environment. The evaluation scripts orchestrated the question bank against the staging orchestrator and generated structured JSON traces. Traces were seamlessly collected via an integrated OpenTelemetry-based observability pipeline, which captured granular telemetry data—including prompt inputs, model outputs, tool invocations, precise execution durations, and the specific agent decisions—directly during runtime without imposing overhead. Each trace recorded system KPIs (e.g., success rate, latency), the LangGraph execution path, and total token cost.
@@ 5.1 RQ1: Effectiveness of the Quality Gate
The framework proved effective at enforcing release discipline. Of 38 evaluation runs, 36 received a PROMOTE decision and 2 were flagged as ROLLBACK (Table 5). The two ROLLBACK decisions occurred during the earliest runs, where evidence coverage dropped to 50%—well below the 56% ROLLBACK threshold (70% of the 80% target). This early detection prevented a build with inadequate citation behavior from being promoted in the staging gate.
@@
Over the four-week evaluation spanning 20+ internal releases, the promoted builds showed no severe regression events in staging gate results, supporting launch-readiness decisions with lower operational risk.
@@ 5.3 RQ2: Predictive Dimensions of Quality
Table 7: Spearman rank correlation matrix between variable quality dimensions (RC excluded as constant).
@@ 6 Discussion
The results of this longitudinal study underscore the necessity of adopting empirical oversight mechanisms like automated self-testing in LLM application environments. Our findings validate that quantitative thresholds mapped to structured conversation paths successfully catch elusive behavioral regressions that traditional testing suites (e.g., unit/integration tests) miss [6].
@@ 6.1 Implications for Industry Practice
The Automated Self-Testing Quality Gates framework represents a practical shift from ad-hoc manual testing to systematic, evidence-driven release management in Agentic AI. The PROMOTE/HOLD/ROLLBACK logic instills confidence in CI/CD pipelines previously rendered untrustworthy by non-deterministic LLM behavior. In this study, persona-based dogfooding was used as a pre-production proxy for customer usage, enabling faster product evolution while preserving quality before first external launch.
Crucially, treating the evaluation dataset as a living, dynamic repository—constantly expanding via active learning and real-world internal testing insights—prevents the “teaching to the test” fallacy. By continuously injecting emerging conversational paradigms into the question bank, organizations can ensure that their quality gates evolve proportionally to reflect the complexity of the underlying LLM models.
Notably, the framework described in this paper was applied to its own development process: an agentic development loop with automated quality gates governed code changes, documentation updates, and manuscript revisions. This meta-application demonstrates practical self-consistency and validates the framework’s utility beyond the specific case study.
@@ 6.3 Evaluator Alignment and Automation-Bias Controls
Table 12: Evaluator alignment study: $n=60$ stratified test cases (20 system-failed, 40 system-passed), evaluated blind by two independent human evaluators and an LLM-as-judge using a different vendor model to reduce shared-bias risk [30]. Human evaluator columns completed. $\kappa$ interpretation follows Landis & Koch [14]; see Section 6 for disaggregated failure-mode analysis.
@@
The LLM-as-judge study reveals complementary coverage between automated structural evaluation and content-focused assessment rather than a simple agreement test. Overall Cohen’s $\kappa$ between the LLM judge and the system gate was $\kappa=0.13$; however, disaggregating by failure type shows this low coefficient is expected and informative. Of the 20 system-rejected cases in the calibration sample, 16 were latency-threshold violations: the response content was substantively acceptable (LLM judge: ACCEPT) but exceeded the 15,000ms P95 threshold, a structural failure invisible in the response text alone. Of the 40 system-accepted cases, 9 were rejected by the LLM judge for content issues: truncated responses, factual fabrication (hallucinated feature names), generic FAQ deflection on specific analytical queries, and language-locale mismatches. This complementary disagreement profile validates the multi-dimensional gate design: structural dimensions (latency, routing, evidence coverage) capture failure classes that content-only evaluation cannot detect, while content assessment exposes false negatives in automated structural evaluation.
@@ 7 Conclusion
For practitioners, the key implication is that release governance for LLM applications should combine content-level assessment with structural runtime signals (e.g., latency, routing, and evidence constraints), rather than relying on either axis alone. For researchers, the next step is external replication across additional domains, model families, and deployment contexts, including direct benchmarked comparisons against alternative automated testing protocols under matched experimental budgets and domain-transfer calibrations as discussed in Section 6.2. Supplementary pseudocode and rubric artifacts are provided in Appendix Section A to support cumulative, externally validated progress in empirical QA for agentic systems.

### 21 May 2026 v2

@@ Abstract
LLM applications are AI systems whose non-deterministic outputs and evolving model behavior make traditional testing insufficient for release governance. We present an automated self-testing framework that introduces quality gates with evidence-based release decisions (PROMOTE/HOLD/ROLLBACK) across five empirically grounded dimensions: task success rate, research context preservation, P95 latency, safety pass rate, and evidence coverage. We evaluate the framework through a longitudinal case study of an internally deployed multi-agent conversational AI system with specific marketing capabilities in active development, covering 38 evaluation runs across 20+ internal releases. The gate identified two ROLLBACK-grade builds in early runs and supported stable quality evolution over a four-week staging lifecycle while exercising persona-grounded, multi-turn, adversarial, and evidence-required scenarios. Statistical analysis (Mann-Kendall trends, Spearman correlations, bootstrap confidence intervals), gate ablation, and overhead scaling indicate that evidence coverage is the primary severe-regression discriminator and that runtime scales predictably with suite size. A human calibration study ($n=60$ stratified cases, two independent evaluators, LLM-as-judge cross-validation) reveals complementary multi-modal coverage: LLM-judge disagreements with the system gate ($\kappa=0.13$) are attributable to structural failure modes—latency violations and routing errors—invisible in response text alone, while the judge independently surfaces content quality failures missed by structural checks, consistent with a multi-dimensional gate design. The framework, supplementary pseudocode, and calibration artifacts are provided to support AI-system quality assurance and independent replication.
@@ 1 Introduction
Releasing LLM-based applications safely is hard because their outputs are probabilistic, prone to “hallucinations,” and heavily context-dependent [5]. Standard unit and integration tests, designed for deterministic software, miss the behavioral regressions that matter most in production: persona drift, subtle safety violations, and unsupported evidence claims. This is particularly visible in domains such as marketing automation and conversational analytics, where multi-agent orchestrators compose tool calls, retrieval, and natural-language reasoning in a single user-visible response.
This mismatch creates a persistent release question for engineering teams: when is a new conversational AI release ready for production? Beyond functional correctness, teams must enforce guardrails consistently and prevent regressions in reasoning behavior. The current ecosystem still lacks automated, data-driven end-to-end testing frameworks that evaluate qualitative dimensions from a realistic user-perception perspective, which limits reliable continuous deployment in LLM systems. Our focus here is narrow: an automated self-testing loop for longitudinal quality-gated deployment of a specific multi-agent application.
To mitigate this gap, we present an empirical testing framework based on Automated Self-Testing (historically known as dogfooding [9]), designed to serve as Quality Gates prior to the release of complex LLM systems. By creating a continuous loop where the system is intensively evaluated against a static and dynamic “question bank,” our framework assesses the build across five empirically grounded dimensions—Task Success Rate (build stability and response utility), Research Context Preservation (multi-turn context continuity), P95 Latency (user experience), Safety Pass Rate (guardrail adherence), and Evidence Coverage (factuality anchoring)—selected to cover both technical robustness (latency, safety) and user-perceived quality (task success, evidence). Based on these metrics, the framework generates a clear release management decision: PROMOTE/HOLD/ROLLBACK.
We validated our method through a longitudinal empirical study (case study) using an internally deployed multi-agent orchestration system (i.e., a conversational AI agent with specific marketing and analytics capabilities) in active development. We analyzed 38 evaluation runs with question banks evolving from 59 to 133 scenarios across more than 20 internal system releases. The framework flagged 2 ROLLBACK-grade builds—both confirmed as severe regressions by the gate ablation analysis (Section 5)—and supported stable quality evolution over a four-week staging lifecycle. Statistical analysis reveals that evidence coverage was the primary determinant of release rejection, while success rate exhibits a controlled decrease as the question bank grows more challenging.
@@
    A formalized and scalable self-testing framework acting as a quality gate (decision checkpoint) for Agentic AI release governance.
@@
    An empirically calibrated set of quality acceptance thresholds for systemic approval (PROMOTE/HOLD/ROLLBACK), accompanied by illustrative pseudocode and sample test cases that enable community adaptation without exposing proprietary assets.
@@ 2.2 Automated Evaluation of LLMs
In contrast to isolated benchmarking, evaluating production systems requires addressing the real-world usage distribution. Our framework builds upon automated evaluation but integrates it directly into the release pipeline as a strict quality gate, rather than an offline benchmarking tool. This emphasis is complementary to broader release-readiness harnesses that combine evaluation, observability, and CI gating across heterogeneous LLM/RAG scenarios [15]. Relative to that broader framing, the present work narrows the scope to automated self-testing as a longitudinal release-control mechanism for a single deployed multi-agent system.
@@ 3.2 Quality Dimensions and Thresholds
  • Latency (P95 <15,000<15{,}000 ms): The 95th percentile response time must not exceed 15 seconds, ensuring acceptable real-time user UX despite multi-agent orchestration overhead.
@@
Task Success Rate ($\geq$80%) was set to permit iterative feature development while blocking builds with systemic failures. Pre-study runs established a natural baseline of approximately 92–95%; the 80% threshold provides a 12–15 percentage-point safety margin, which proved sufficient to absorb the observed suite-size expansion (Section 5) without spurious rollbacks in the 38-run window.
@@
P95 Latency (<15,000<15{,}000 ms) reflects the upper bound of acceptable wait time for an interactive marketing analytics assistant. The threshold was set operationally from pre-launch internal testing, where responses exceeding 15 seconds consistently triggered user abandonment in dogfooding sessions; it is loose relative to traditional web-UX response-time guidance but accommodates the multi-agent orchestration overhead inherent to the system under test.
@@ 3.4 Decision Engine
This evidence-driven approach ensures that behavioral regressions and subtle degradations—such as an increase in hallucination rates disguised as confident answers—are caught deterministically before impacting users. The deterministic decision logic is given in Algorithm 1, and the dimensional metric extraction in Algorithm 2 (Appendix A). Figure 3 summarizes the release-decision state machine, and Table 4 positions automated self-testing relative to traditional U/I and E2E testing coverage.
@@ 4.3 Methodology and Data Collection
During this longitudinal period, every code merge targeting the main branch triggered an automated self-testing workflow within an isolated staging environment. The evaluation scripts orchestrated the question bank against the staging orchestrator and generated structured JSON traces. Traces were seamlessly collected via an integrated OpenTelemetry-based observability pipeline, which captured granular telemetry data—including prompt inputs, model outputs, tool invocations, precise execution durations, and the specific agent decisions—directly during runtime with negligible runtime overhead. Each trace recorded system KPIs (e.g., success rate, latency), the LangGraph execution path, and total token cost.
@@ 5.1 RQ1: Effectiveness of the Quality Gate
The framework proved effective at enforcing release discipline. Of 38 evaluation runs, 36 received a PROMOTE decision and 2 were flagged as ROLLBACK (Table 5). The two ROLLBACK decisions occurred during Run 1 and Run 2 (the first two evaluations in the longitudinal series), where evidence coverage dropped to 50%—well below the 56% ROLLBACK threshold (70% of the 80% target). This early detection prevented a build with inadequate citation behavior from being promoted in the staging gate.
@@
Over the four-week evaluation spanning 20+ internal releases, no promoted build subsequently triggered a rollback in later runs, indicating stable post-promote behavior under the gate’s criteria.
@@ 5.3 RQ2: Predictive Dimensions of Quality
Table 7: Spearman rank correlation matrix between variable quality dimensions (RC excluded as constant). $p<0.05$ (two-sided, $n=38$); coefficients without a marker are not statistically significant.
@@ 6 Discussion
The results of this longitudinal study support the value of adopting empirical oversight mechanisms like automated self-testing in LLM application environments. Our findings show that quantitative thresholds mapped to structured conversation paths can catch behavioral regressions that traditional testing suites (e.g., unit/integration tests) often miss [5].
@@ 6.1 Implications for Industry Practice
The Automated Self-Testing Quality Gates framework represents a practical shift from ad-hoc manual testing to systematic, evidence-driven release management in Agentic AI. The PROMOTE/HOLD/ROLLBACK logic instills confidence in CI/CD pipelines previously rendered unreliable by non-deterministic LLM behavior. In this study, persona-based dogfooding was used as a pre-production proxy for customer usage, enabling faster product evolution while preserving quality before first external launch.
Crucially, treating the evaluation dataset as a living, dynamic repository—constantly expanding via active learning and real-world internal testing insights—mitigates overfitting to the evaluation set. By continuously injecting emerging conversational paradigms into the question bank, organizations can ensure that their quality gates evolve proportionally to reflect the complexity of the underlying LLM models.
@@ 6.3 Evaluator Alignment and Automation-Bias Controls
Table 12: Evaluator alignment study: $n=60$ stratified test cases (20 system-failed, 40 system-passed), evaluated blind by two independent human evaluators and an LLM-as-judge using a judge model from a different capability tier within the same model family (partial mitigation of shared-bias risk; same-family bias cannot be fully excluded) [31]. Human evaluator columns completed. $\kappa$ interpretation follows Landis & Koch [13]; see Section 6 for disaggregated failure-mode analysis.
@@
The LLM-as-judge study reveals complementary coverage between automated structural evaluation and content-focused assessment rather than a simple agreement test. Overall Cohen’s $\kappa$ between the LLM judge and the system gate was $\kappa=0.13$; however, disaggregating by failure type shows this low coefficient is expected and informative. The 20 system-rejected cases split into 16 latency-threshold violations and 4 routing/metric-extraction failures. The LLM judge accepted 13 of the 20 system rejections (9 of the 16 latency-only rejects, plus all 4 routing failures), indicating that for these cases the response content was substantively acceptable but the build failed on structural dimensions (P95 latency above the 15,000 ms threshold or an incorrect agent path) that are invisible in the response text alone. Conversely, of the 40 system-accepted cases, 9 were rejected by the LLM judge for content issues: truncated responses, factual fabrication (hallucinated feature names), generic FAQ deflection on specific analytical queries, and language-locale mismatches. This complementary disagreement profile is consistent with a multi-dimensional gate design: structural dimensions (latency, routing, evidence coverage) capture failure classes that content-only evaluation may miss, while content assessment exposes false negatives in automated structural evaluation.
@@ 7 Conclusion
For practitioners, the key implication is that release governance for LLM applications should combine content-level assessment with structural runtime signals (e.g., latency, routing, and evidence constraints), rather than relying on either axis alone. For researchers, the next step is external replication across additional domains, model families, and deployment contexts, including direct benchmarked comparisons against alternative automated testing protocols under matched experimental budgets and domain-transfer calibrations as discussed in Section 6.2. Supplementary pseudocode and rubric artifacts are provided in Appendix Section A to support cumulative, externally validated progress in empirical QA for agentic systems. More broadly, we view automated self-testing as one specialization of a larger release-readiness stack in which observability, reproducible artifacts, and CI gating are shared across projects and evaluation modes [15].